1. If a consumer increases her quantity of ice cream consumed by 100% when her

income rises by 25%. Calculate her income elasticity of demand for the ice cream and

interpret the result. (2marks)


2. If Qs = -20 + 10p, and Qd = 400 - 20p, what is the equilibrium price and quantity? (2

marks)


3. Assume you are managing a food processing plant in Ethiopia. The demand function

for one of your product is given as Qd=50-2p. (8 marks)

a) Find the point price elasticity if price is 15 ETB? Is it elastic or inelastic?

b) How do you interpret the elasticity result?

c) In order to get more revenue what will be your recommendation. Is it to increase

price or decrease price? Why?

d) Describe at least four determinants of the price elasticity demand for the food

product?


1
Expert's answer
2021-06-14T13:34:31-0400

Solution

a.) Elasticity of demand="\frac {100}{25}" = 4


The demand elasticity for the ice cream is elastic since it's greater than 1.


2.Qs=-20+10p

Qd=400-20p

Qs=Qd

400-20p=-20+10p

420=30p

P=14

Q=120


3a.)PED="\frac{P}{Q}.\frac{∆Q}{∆P}"


"\frac{∆Q}{∆P}=\frac{-2}{1}=-2"


Qd=50-2(15)=20

PED="\frac{15}{20}\times(-2)=-1.5"


PED is inelastic


b.) The point price elasticity is less than 1 rather negative making it inelastic


c.) For price inelastic goods or services, the change in price will have minimal effect to the total revenue this is because the quantity demanded is not affected by change in price of goods.


d.)

  • Substitutes availability
  • Period time taken since price changed
  • Amount of income spent on the good
  • If the good is a necessity or luxury